OpenStudy (anonymous):

Solve the equation below by completing the square and then solving for ('x') x^2+x+1=7/4 HELP !!! Please show steps ! Thank you ! A. x=1 or x=-1 B. x=1/2 or x=-1/2 C. x=-3/2 or x=-1/2 D.x =1 or x=1

OpenStudy (anonymous):

x^2+x+1=7/4 -> (x^2+x+1) - 1 = 7/4 - 1 -> x^2+x = 3/4 -> x^2+x + (1/2)^2 = 3/4 + (1/2)^2 -> (x + 1/2)^2 = 1. Take square roots. x + 1/2 = +-1 -> x = -3/2 or -1/2

OpenStudy (anonymous):

Thank you !